# Noncommutative Minkowski integral inequality and a unitary categorification criterion for fusion rings

### v1

@@ 1 Introduction
We tested the above inequality on the aforementioned dataset of fusion rings from [VS23]. Out of 28,450 fusion rings, 5,793 were excluded, corresponding to 20.36%. Among them, 540 items, that is, 1.898%, cannot be excluded by the primary $3$-criterion of [HLP+24]. Moreover, $23$ of these cannot be excluded by the $d$-number, extended cyclotomic, Lagrange, and pseudo-unitarity Drinfeld criteria in [LPR22]. Some of these cannot even be excluded by the criterion in [ABD+25, Remark 3.2].
@@ 2.1 Reduced Hölder inequality
The equality in (6) must also hold. From (5) it follows that the maximum modulus $1$ of $f(w)\left[f(1)^{w}f(0)^{1-w}\right]^{-1}$ is attained at the interior point $1/p$ of $\Omega$. Thus, by the maximum modulus principle,
@@ 3.1 Construction of commuting squares
of multi-matrix algebras. Let $m_{ij}$ be the multiplicity of the summand $M_{k_{i}}(\mathbb{C})$ in $M_{l_{j}}(\mathbb{C})$. Then the matrix $T=(m_{ij})_{i,j}$ is called an inclusion matrix. Let $\operatorname{tr}:A_{1}\rightarrow\mathbb{C}$ be a faithful tracial state on $A_{1}$ and let $v_{j}=\operatorname{tr}(q_{j})$ where $q_{j}$ is a minimal projection in the direct summand $M_{l_{j}}(\mathbb{C})$ of $A_{1}$. Then the (column) vector $v=(v_{j})_{j}$ is called the trace vector corresponding to $\operatorname{tr}$. It follows that $Tv$ is the trace vector corresponding to the induced trace $\operatorname{tr}|_{A_{0}}$ of $A_{0}$.
@@ 3.2 Unitary categorification of fusion rings
is a commuting square of multi-matrix algebras where the conditional expectations are the normalized partial traces, and hence, we can apply Theorem 3.2. Replacing $(T_{10}^{11})_{**}$, $(T_{00}^{10})_{1*}$, $(T_{01}^{11})_{**}$ and $(T_{00}^{01})_{1*}$ in (14) by $N_{*k}^{*}$, $N_{ij}^{*}$, $N_{i*}^{*}$ and $N_{jk}^{*}$, and replacing $v_{*}^{11}$, $v_{*}^{10}$, $v_{*}^{01}$ and $v_{*}^{00}$ by $\frac{d_{*}}{d_{i}d_{j}d_{k}}$, $\frac{d_{*}}{d_{i}d_{j}}$, $\frac{d_{*}}{d_{j}d_{k}}$ and $1$, respectively, gives
@@
      There exists an involution $i\mapsto\overline{i}$ of $I$ such that the induced map $\sum_{i\in I}n_{i}b_{i}\mapsto\sum_{i\in I}n_{i}b_{\overline{i}}$ is an anti-isomorphism of rings;
@@
Grothendieck rings of fusion categories are fusion rings. Corollary 3.4 can be used to disqualify unital based rings from having unitary categorifications: If there is a triple of basis elements $c_{i},c_{j},c_{k}$ of a unital based ring $\mathcal{R}$, nonnegative numbers $(a_{n})_{n}$ and $1\leq p<\infty$ that do not satisfy Corollary 3.4, then $\mathcal{R}$ does not admit a unitary categorification. In particular, this criterion can be applied to fusion rings. However, difficulties lie in choosing the appropriate $(a_{n})_{n}$ and $p$. We applied the following algorithm to each triple $(c_{i},c_{j},c_{k})$ for our calculations:
@@
The numbers $a_{n}$ and $1/p$ are sampled from the uniform distribution. To approximate the partial derivative of $\operatorname{RHS}/\operatorname{LHS}$ with respect to $a_{n}$ and $1/p$, we computed the difference quotient with $\Delta a_{n}=\Delta(1/p)=0.0001$. The gradient descent method is applied with step size $0.05$ and $1,000$ iterations. The overall process is implemented in MATLAB and is repeated up to $20$ times for each choice of a fusion ring.
We tested the algorithm on the dataset of $28,450$ fusion rings from [VS23] and excluded $5,793$ of them (20.36%). Among them, $540$ fusion rings ($1.898\%$) cannot be excluded by the primary $3$-criterion [HLP+24]. Moreover, $23$ of these cannot be excluded by the following criteria in [LPR22]: $d$-number, extended cyclotomic, Lagrange, and pseudo-unitarity Drinfeld. Some of these admit induction matrices and thus cannot be excluded by the criterion in [ABD+25, Remark 3.2]. For more details of the results, see Tables 2 and 2. The author thanks Palcoux for testing the criteria in [LPR22] and [ABD+25].
@@
The following fusion rings pass the primary $3$-criterion, but fail to satisfy (15). Here, the $(n,m)$-entry of the $l$-th matrix is $N_{lm}^{n}$.
@@
    The following is the simplest such a fusion ring: [1 0000 1000 0100 001 ][0 1001 1110 1000 100 ][0 0010 1001 0000 010 ][0 0100 1000 0011 000 ] For $i=j=k=2$, $\frac{1}{p}=0.661975038859587$, and
@@
    The following is the simplest among those that do not even pass the criteria in [LPR22] and [ABD+25]:

### v2

@@ 1 Introduction
We tested the above inequality on the aforementioned dataset of fusion rings from [VS23]. Out of 28,450 fusion rings, 5,793 were excluded, corresponding to 20.36%. Among them, 540 items, that is, 1.898%, cannot be excluded by the primary 33-criterion of [HLP+24]. Moreover, $23$ of these cannot be excluded by the $d$-number, extended cyclotomic, Lagrange, and pseudo-unitarity Drinfeld criteria in [LPR22]. Some of these cannot even be excluded by the criterion in [ABD+25, Remark 3.2].
@@ 2.1 Reduced Hölder inequality
The equality in (6) must also hold. From (5) it follows that the maximum modulus 11 of $f(w)\left[f(1)^{w}f(0)^{1-w}\right]^{-1}$ is attained at the interior point $1/p$ of $\Omega$. Thus, by the maximum modulus principle,
@@ 3.1 Construction of commuting squares
of multi-matrix algebras. Let $m_{ij}$ be the multiplicity of the summand $M_{k_{i}}(\mathbb{C})$ in $M_{l_{j}}(\mathbb{C})$. Then the matrix $T=(m_{ij})_{i,j}$ is called an inclusion matrix. Let $\operatorname{tr}:A_{1}\rightarrow\mathbb{C}$ be a faithful tracial state on $A_{1}$ and let $v_{j}=\operatorname{tr}(q_{j})$ where $q_{j}$ is a minimal projection in the direct summand $M_{l_{j}}(\mathbb{C})$ of $A_{1}$. Then the (column) vector $v=(v_{j})_{j}$ is called the trace vector corresponding to $\operatorname{tr}$. It follows that TvTv is the trace vector corresponding to the induced trace $\operatorname{tr}|_{A_{0}}$ of $A_{0}$.
@@ 3.2 Unitary categorification of fusion rings
is a commuting square of multi-matrix algebras where the conditional expectations are the normalized partial traces, and hence, we can apply Theorem 3.2. Replacing $(T_{10}^{11})_{**}$, $(T_{00}^{10})_{1*}$, $(T_{01}^{11})_{**}$ and $(T_{00}^{01})_{1*}$ in (14) by $N_{*k}^{*}$, $N_{ij}^{*}$, $N_{i*}^{*}$ and $N_{jk}^{*}$, and replacing $v_{*}^{11}$, $v_{*}^{10}$, $v_{*}^{01}$ and $v_{*}^{00}$ by $\frac{d_{*}}{d_{i}d_{j}d_{k}}$, $\frac{d_{*}}{d_{i}d_{j}}$, $\frac{d_{*}}{d_{j}d_{k}}$ and 11, respectively, gives
@@
      There exists an involution $i\mapsto\overline{i}$ of II such that the induced map $\sum_{i\in I}n_{i}b_{i}\mapsto\sum_{i\in I}n_{i}b_{\overline{i}}$ is an anti-isomorphism of rings;
@@
Grothendieck rings of fusion categories are fusion rings. Corollary 3.4 can be used to disqualify unital based rings from having unitary categorifications: If there is a triple of basis elements $c_{i},c_{j},c_{k}$ of a unital based ring $\mathcal{R}$, nonnegative numbers $(a_{n})_{n}$ and $1\leq p<\infty$ that do not satisfy Corollary 3.4, then $\mathcal{R}$ does not admit a unitary categorification. In particular, this criterion can be applied to fusion rings. However, difficulties lie in choosing the appropriate $(a_{n})_{n}$ and pp. We applied the following algorithm to each triple $(c_{i},c_{j},c_{k})$ for our calculations:
@@
The numbers $a_{n}$ and $1/p$ are sampled from the uniform distribution. To approximate the partial derivative of $\operatorname{RHS}/\operatorname{LHS}$ with respect to $a_{n}$ and $1/p$, we computed the difference quotient with $\Delta a_{n}=\Delta(1/p)=0.0001$. The gradient descent method is applied with step size $0.05$ and $1,000$ iterations. The overall process is implemented in MATLAB and is repeated up to 2020 times for each choice of a fusion ring.
We tested the algorithm on the dataset of $28,450$ fusion rings from [VS23] and excluded $5,793$ of them (20.36%). Among them, $540$ fusion rings ($1.898\%$) cannot be excluded by the primary 33-criterion [HLP+24]. Moreover, $23$ of these cannot be excluded by the following criteria in [LPR22]: $d$-number, extended cyclotomic, Lagrange, and pseudo-unitarity Drinfeld. Some of these admit induction matrices and thus cannot be excluded by the criterion in [ABD+25, Remark 3.2]. For more details of the results, see Tables 2 and 2. The author thanks Palcoux for testing the criteria in [LPR22] and [ABD+25].
@@
The following fusion rings pass the primary 33-criterion, but fail to satisfy (15). Here, the $(n,m)$-entry of the $l$-th matrix is $N_{lm}^{n}$.
@@
    The following is the simplest such a fusion ring: [1 $00$ 1000 0100 001 ][0 1001 1110 1000 100 ][0 0010 1001 $00$ 010 ][0 0100 1000 0011 000 ] For $i=j=k=2$, $\frac{1}{p}=0.661975038859587$, and
@@
    The following is the simplest integral example among those that cannot be even ruled out by the criteria in [LPR22] and [ABD+25]: